Analysis of the water content of a lake found in the desert showed that it contained 17.5 percent chloride ion, and had a density of 1.23 g/mL. What is the molarity of the chloride ion in the water?
Answer #1

TU Molarity (M) = (W/W) 1. + d +10 GMW = 17.5+ 1.23710 35.5 molarity = 6.06 M